OpenStudy (anonymous):

Angular momentum of 3p orbital

OpenStudy (anonymous):

@mathslover

OpenStudy (anonymous):



mathslover (mathslover):

l = ? (u know the formula? )

mathslover (mathslover):

\(\sqrt{l(l+1) } \times \cfrac{h}{2\pi}\) what is l for 3p orbital?

OpenStudy (anonymous):

Is it root

mathslover (mathslover):

yep.. orbital angular momentum = \(\sqrt{l(l+1)} \cfrac{h}{2\pi}\)

OpenStudy (anonymous):

O_o Ncert mw without root likha hai S orbital ke liye to without root kuch farak nahi padta so i got the first answer But for p i got 1/pi according to ncert "so called formula" abhi answer milgaya

OpenStudy (anonymous):

Ncert has without root lol

mathslover (mathslover):

:-) koi nahin, answer aa gaya na? bass toh fir party karte hain.. :P

OpenStudy (anonymous):

angular momentum of 4th bohr orbit?

mathslover (mathslover):

hmm, try kiya?

OpenStudy (anonymous):

milgaya sorry

mathslover (mathslover):

:)

OpenStudy (anonymous):

wait is it 2h/pi

mathslover (mathslover):

l kya hota hai 4th bohr orbit ke liye?

OpenStudy (anonymous):

I used mvr=nh/2pi

mathslover (mathslover):

okay, so u just put n = 4 there, right?

OpenStudy (anonymous):

yeah

mathslover (mathslover):

seems right to me.

OpenStudy (anonymous):

you are attending the plancess ........

mathslover (mathslover):

yeah

OpenStudy (anonymous):

okay